OpenStudy (anonymous):

Hugo and Viviana work in an office with ten other coworkers. Out of these 12 workers, their boss needs to choose a group of five to work together on a project. Suppose Hugo and Viviana absolutely refuse, under any circumstances, to work together. Under this restriction, how many different working groups of five can be formed?

OpenStudy (anonymous):

Its Discrete Mathematics fyi DemolisionWolf

OpenStudy (anonymous):

Hugo and Viviana work in an office with ten other coworkers. 12 workers, (thus n=12) their boss needs to choose a group of five (thus r=5) Suppose Hugo and Viviana absolutely refuse, under any circumstances, to work together. (this line, im still thinking about.) But we want to use the 'combination probability' formula

OpenStudy (anonymous):

whats your idea on this part? "Suppose Hugo and Viviana absolutely refuse, under any circumstances, to work together." up to this point, i think n=12 and r = 5, but this phrase is messing me up..

OpenStudy (anonymous):

Actually what I'm thinking is that 12-2 = 10, so 12^10 + 5^10

OpenStudy (anonymous):

Unfortunately, my homework still says it wrong

OpenStudy (anonymous):

have you used the combination formula? like... n!/ (n-r)!

OpenStudy (anonymous):

I just used recursive math to assume my solution, I think I'm either lost or on the wrong track, I will try to plug in those numbers to that equation & see if it works tho

OpenStudy (anonymous):

did it work?

OpenStudy (anonymous):

oh maybe it's n=12-1=11 so we get rid of one of the two bosses but not both?

OpenStudy (anonymous):

I did 12!/(12-5) = 479001600/7 = 68428800 which is wrong

OpenStudy (anonymous):

I can try it with 11, hold on

OpenStudy (anonymous):

try this one, I think I had the combination wrong before. and do n=11, r=5 n!/(r(n-r)!)

OpenStudy (anonymous):

11! = 39916800/30 = 1330560 but still incorrect, I have a feeling since Hugo & vivian cannot work together, I will try 10!

OpenStudy (anonymous):

don't for get the ! for the denominator term

OpenStudy (anonymous):

Omg completely missed that, in that case my calculation was incorrect, hold on

OpenStudy (anonymous):

11! = 39916800/2.652528598121911e+32 = 1.50485842e-39

OpenStudy (anonymous):

Not correct tho

OpenStudy (anonymous):

hmm.. sorry i'm not sure what else to try :(

OpenStudy (anonymous):

Np I really appreciate all your help man